OpenStudy (anonymous):

Show that the given function is a pdf on the indicated interval. f(x) = 4x3, [0, 1]

OpenStudy (dan815):

pdf?

OpenStudy (anonymous):

Probability Density Functions

OpenStudy (anonymous):

@Luigi0210 can u help?

OpenStudy (luigi0210):

@primeralph @dan815 @johnweldon1993

OpenStudy (anonymous):

@abb0t

OpenStudy (abb0t):

Well, if you notice, with ur points it's outside of the given range 0<x<1 so to show that it's a probability density, you gotta show \(\large \int\limits_{-\infty}^{\infty}f(x)dx =1\)

OpenStudy (abb0t):

remember that f(x) must be >0 for all x. that was another one of the conditions

OpenStudy (abb0t):

That's a fairly simple function to integrate so i'll leave it to you to integrate.

OpenStudy (anonymous):

okay thanks

OpenStudy (abb0t):

You're very welcome. If you're still confused, don't hesitate to ask.

OpenStudy (anonymous):

ok